If f (x) = 2x^2 − 3x + 1, then f (x −3) = ? Express your answer as a polynomial in x.
Answer by stanbon(75887)

If f (x) = 2x^2 − 3x + 1, then f (x −3) = ?
Express your answer as a polynomial in x.
-----------------------
f(x-3) = 2(x-3)^2 - 3(x-3) + 1
---
f(x-3) = 2[x^2-6x+9] -3x + 9 + 1
---
f(x-3) = 2x^2-12x+18-3x+10
---
f(x-3) = 2x^2-15x+28
========================
Cheers,
Stan H.